Find the probability that x falls in the shaded area. (Enter an exact number as an integer, fraction, or decimal.) 0 12 3 45 6 7 8 9 10
Answer #1

Solution :

Given that,

1 / (b - a) = 1 / 9

P(c < x < d) = (d - c) / (b - a)

P(2 < x < 3) = (3 - 2) / 9 = 1 / 9

Probability = 1 / 9